# MMSD3.0: A Multi-Image Benchmark for Real-World Multimodal Sarcasm Detection

### v1

@@ Abstract
Despite progress in multimodal sarcasm detection, existing datasets and methods predominantly focus on single-image scenarios, overlooking potential semantic and affective relations across multiple images. This leaves a gap in modeling cases where sarcasm is triggered by multi-image cues in real-world settings. To bridge this gap, we introduce MMSD3.0, a new benchmark composed entirely of multi-image samples curated from tweets and Amazon reviews. We further propose the Cross-Image Reasoning Model (CIRM), which performs targeted cross-image sequence modeling to capture latent inter-image connections. In addition, we introduce a relevance-guided, fine-grained cross-modal fusion mechanism based on text–image correspondence to reduce information loss during integration. We establish a comprehensive suite of strong and representative baselines and conduct extensive experiments, showing that MMSD3.0 is an effective and reliable benchmark that better reflects real-world conditions. Moreover, CIRM demonstrates state-of-the-art performance across MMSD, MMSD2.0 and MMSD3.0, validating its effectiveness in both single-image and multi-image scenarios.
@@ 1 Introduction
Sarcasm is a common affective expression, yet it is hard to capture because the literal wording often contradicts the speaker’s true attitude and it appears in diverse ironic forms. Accurately identifying sarcasm is therefore of substantial importance for emotion detection, online services, and opinion mining [36, 46, 22, 44, 49].
With the rapid surge of multimodal content on social media platforms, the expression of sarcasm has become increasingly diverse. In particular, sarcasm conveyed through both images and text poses significant challenges for automatic detection. This has led to the emergence of the multimodal sarcasm detection task. Cai et al. [3] were the first to construct the widely used MMSD dataset based on Twitter, providing a benchmark for research in this area. Early approaches [3, 44, 23, 41, 31] built on MMSD fused visual and textual modalities to capture cross-modal incongruity and have achieved encouraging results. However, Qin et al. [32] later identified the presence of spurious cues and problematic samples within MMSD, and subsequently introduced an improved version, MMSD2.0. This new dataset addressed these issues and further advanced the development of more reliable multimodal sarcasm detection. [37]
@@
According to Zhang et al. [48], over 42% of tweets contain more than one image, which is a notably high proportion. In such multi-image scenarios, sarcasm often arises from latent semantic or affective associations among the images. As shown in Figure 1, sarcasm can emerge from the contrast between images, for instance, the left image depicts Laura Loomer, while the right shows Count von Count. Without both images, the key contrast driving the sarcastic intent cannot be perceived. However, existing datasets [3, 32] and methods [3, 44, 30, 23, 41, 31, 32, 40, 4, 37] focus solely on single-image settings and cannot capture sarcasm driven by relations across multiple images. As a result, they do not faithfully reflect the complexity of multimodal sarcasm in real-world contexts.
@@
At the same time, to address the challenges of multi-image sarcasm detection, we propose CIRM. CIRM is a novel architecture for cross-image and cross-modal reasoning. It encodes image order through positional embeddings and masking, and employs a Dual-Stage Bridge Module to enable interaction before and after sequence modeling. Furthermore, a relevance-guided fine-grained fusion module leverages OCR-based alignment and relevance weights to selectively aggregate visual–textual cues into a unified representation.
@@
  • We propose CIRM, a multi-image sarcasm reasoning framework that performs dual-stage cross-modal bridging and relevance-guided fusion to enhance fine-grained text–image alignment and improve multimodal understanding in complex sarcastic scenarios.
@@ 2.2 Multimodal Large Language Models
Table 1: Statistics of MMSD3.0 dataset across different splits.
Split Total Sar. Non-sar. Sar. (%) Avg. Len / Img
Train 7583 3007 4576 39.65 31.81 / 2.57
@@ 3.1 Data Collection
AI-generated Content and Augmentation. To mirror the surge of AI-generated deception on Twitter [35, 5, 10] and Amazon [45, 26, 29], we augment our data with Qwen2.5-VL-32B [1] as the generator and GPT-4o as the evaluator. With two platform-specific prompts 8 and  9, we feed the images from 1,444 real samples to produce three sarcastic candidates per sample. As shown in Appendix Figure 10, the evaluation prompt then rates candidates on six criteria (e.g., Naturalness, Authenticity) and selects the top-scoring one, yielding 1,444 sarcasm-labeled samples.
@@ 3.2 Human Annotation
We recruited nine graduate students with backgrounds in multimodal research to conduct manual annotation. Given the inherently subjective nature of sarcasm [11], each sample was assigned to two different annotators to reduce potential bias. Since understanding sarcasm in some cases may require specific background knowledge, annotators were allowed to search online for relevant context to better interpret the intended meaning. During the annotation process, we excluded any content containing violence, pornography, or other harmful or sensitive material. To evaluate annotation consistency, we used Cohen’s Kappa [28], and obtained a score of 0.816, indicating a high level of agreement [21]. Finally, we split the dataset into training, validation, and test sets following a 70:15:15 ratio.
@@ 4.4 Dual-Stage Bridge Module
Post-Bridge and Final Representation. After independent sequence modeling within each modality, we perform a second bridging stage to re-establish cross-modal alignment and enhance joint semantics for classification. The sequential modeling block is first applied to both textual and visual streams:
@@
4.5 Relevance-Guided Fusion
@@
Relevance Estimation. A global textual summary $\bar{t}$ is derived by averaging $T^{O}$. Each image representation $v_{i}^{O}$ is then scored against $\bar{t}$ from two perspectives: a cosine similarity term $s_{i}^{\cos}$ and a learnable term $s_{i}^{\mathrm{lrn}}$:
$s_{i}^{\cos}=\cos(W_{v}v_{i}^{O},\,W_{t}\bar{t}),\qquad s_{i}^{\mathrm{lrn}}=\mathrm{MLP}\!\left([v_{i}^{O};\,\bar{t}]\right)$ (17)
@@ 4.6 Classification
After the Dual-Stage Bridge and Relevance-Guided Fusion, we obtain three complementary representations: the post-bridge text and vision streams $(T_{\text{post}},V_{\text{post}})$ and the relevance-guided feature $f$. We aggregate and integrate them in a shared fusion space for final prediction.
@@
where $c_{i}\in\{0,1\}$ masks padded images. The pooled features $\bar{t}$ and $\bar{v}$ are concatenated with the relevance-guided vector $f$ and an optional star-rating embedding:
$z=\mathrm{MLP}_{\text{fuse}}\!\left([\bar{t};\,\bar{v};\,f;\,\mathrm{Emb}(\text{rating})]\right)$ (21)
@@
$\mathcal{L}=\mathrm{CrossEntropy}(\hat{\mathbf{y}},\mathbf{y};\mathbf{w})$ (23)
@@ 5 Experiments
Appendix. See Appendix 5.8 for the non-tiling multi-image protocol (adapting single-image baselines for multiple images), hyperparameter sweeps 5.9, and error cases 5.11.
@@ 5.1 Datasets
Because our data extraction differs from MMSD and MMSD2.0, we do not merge them. We first evaluated CIRM’s single-image performance on the two public datasets, MMSD and MMSD2.0. MMSD2.0 is an improved version of MMSD, with spurious cues removed and annotations refined, using training, validation, and test splits of 80%, 10%, and 10%, respectively. We then constructed a baseline on MMSD3.0 to evaluate multi-image performance, using training, validation, and test splits of 70%, 15%, and 15%.
@@ 5.2 Experimental Settings
We use Qwen2.5-VL-32B [1] as the generator and GPT-4o [16] as the evaluator. During the baseline implementation, we reproduce each method following its official hyperparameters. In CIRM, we use AdamW [27] as the optimizer. The learning rate is set to $2e^{-5}$, with a weight decay of $1e^{-5}$. The batch size is set to 16, and the model is trained for 30 epochs. All baseline construction and model training are conducted on a single NVIDIA H100 GPU with 80GB of memory.
@@ 5.3 Baselines
Setting. We evaluate CIRM in both single-image and multi-image scenarios. For methods that only accept one image, we tile all $N$ images in order into a single canvas and keep the text unchanged, ensuring identical visual evidence while keeping model designs unchanged.
Single-image. CIRM is evaluated against unimodal and multimodal baselines. Text-only: TextCNN [20] (CNN over $n$-grams); Bi-LSTM [12] (bidirectional RNN for long-range context); SMSD [43] (BiLSTM with self-matching, low-rank pooling). Image-only: ResNet [14] (residual CNN); ViT [9] (Transformer over patch tokens). Multimodal: HFM [3] (hierarchical fusion of image, text, attributes); Att-BERT [30] (inter-modal attention on BERT); HKE [23] (graph reasoning over compositional congruity); DIP [41] (factual + affective cues mining); Multi-view CLIP [32] (diverse CLIP views for incongruity); Multi-view CLIP+TFCD [50] (training-free counterfactual debiasing); MoBA [42] (bi-directional LoRA with MoE routing); G2SAM [40] (graph-based global perception); CofiPara [4] (LLM-assisted sarcasm reasoning); Tang et al. [37] (instruction templates with demo retrieval); RCLMuFN [39] (relational learning with multiplex fusion).
Multi-image. CIRM and modern MLLMs natively accept multiple images; for single-image methods we use the tiling protocol above. Text-only: Bi-LSTM [12] (bidirectional recurrent backbone); CLIP-Text [33] (contrastively pre-trained text encoder); BERT [8] (masked-language Transformer baseline); RoBERTa (TweetEval) [2] (robustly trained BERT variant for social text). Image-only: ResNet [14] (residual CNN on composite canvas); ViT [9] (patch-token Transformer on tiled images). Multimodal: DIP [41] (joint factual/affective fusion); Multi-view CLIP [32] (multi-perspective CLIP alignment); MoBA [42] (bidirectional modality interaction for robust fusion); Tang et al. [37] (LLM prompting with retrieved demonstrations). MLLMs: GPT-4o [16] (end-to-end multimodal reasoning with native multi-image input); LLaVA-1.5-7B [24] (vision-language fine-tuning on conversational data); Qwen2.5-VL-32B [1] (large-scale VL model with strong perception and instruction following). Prompts used for these MLLMs are provided in Appendix figure 11.
@@ 5.5 Ablation Study
To verify the contribution of each component, we conduct ablation experiments under several settings: (1) w/o DSBM, removing the Dual-Stage Bridge Module; (2) w/o PE, removing positional encoding and masking; (3) w/o OCR, excluding OCR-based textual cues; (4) w/o RGF, removing the Relevance-Guided Fusion; and (5) removing specific submodules within the Dual-Stage Bridge, including Pre-Bridge, Sequential Modeling, and Post-Bridge.
Table 4 summarizes the results. Removing the full DSBM or the relevance-guided fusion module causes the largest decline, highlighting their crucial roles in capturing cross-modal dependencies. Excluding positional embeddings also weakens the model, showing that encoding image order benefits multi-image understanding. Omitting the Pre-Bridge or Post-Bridge slightly lowers performance, indicating that alignment before and after sequential fusion contributes meaningfully. The absence of OCR causes a clear drop, as many images include textual elements that visual features alone cannot represent. Overall, these findings confirm that each module jointly strengthens the model’s multimodal reasoning and robustness on MMSD3.0.
Modality Method Real-world AI-Gen
Acc (%) F1 (%) Acc (%)
Multimodal DIP  [41] 79.59 75.50 97.98
Multi-view CLIP  [32] 80.01 75.48 95.96
MoBA  [42] 76.02 68.80 88.89
Tang et al.  [37] 80.36 75.93 95.45
MLLM GPT-4o  [16] 71.34 67.60 82.61
LLaVA-1.5-7B  [24] 59.62 55.78 72.83
Qwen2.5-VL-32B  [1] 68.90 66.65 95.65
Ours CIRM 83.31 80.39 98.48
5.6 Real-World Situation
As shown in Table 5, model performance on real-world data remains noticeably lower than on AI-generated samples, reflecting the greater ambiguity and contextual variability of authentic posts. Among multimodal baselines, DIP and Tang et al. perform relatively well, while MoBA lags behind and Multi-view CLIP shows limited generalization. Large multimodal models such as GPT-4o and Qwen2.5-VL-32B achieve moderate accuracy, whereas LLaVA-1.5-7B performs the weakest. CIRM attains the highest F1 score on both data types, reaching $80.39\%$ on real-world and $98.48\%$ on AI-generated samples. These results highlight its robustness across domains and verify that the proposed bridging and relevance-guided fusion effectively enhance visual–textual reasoning, allowing the model to better interpret the nuanced and diverse nature of real multimodal sarcasm.
@@ 5.8 No-Tiling Multi-Image Experiment
Figure 6: Performance variation with mixing weight $\alpha$ on MMSD3.0. The model achieves the best F1-score of $84.42\%$ at $\alpha=0.3$, showing stable results for moderate values.
@@
5.9 Hyperparameter Study on Mixing Weight $\alpha$
We investigate the influence of the mixing weight $\alpha$ in the relevance module, where $s_{i}=\alpha\,s_{i}^{\mathrm{cos}}+(1-\alpha)\,s_{i}^{\mathrm{lrn}}$ balances cosine alignment with the learned scorer. We sweep $\alpha\in\{0,0.1,\ldots,1.0\}$ while keeping other settings fixed and evaluate on the MMSD3.0 validation split. As shown in Figure 6, performance peaks at $\alpha=0.3$, achieving the highest F1-score of $84.42\%$. The accuracy remains stable around $84\%$ across most values, indicating robustness to moderate changes. Both extremely small and large $\alpha$ values, however, slightly reduce performance; for instance, F1 is $83.98\%$ at $\alpha=0.0$ and $83.63\%$ at $\alpha=1.0$. These results suggest that an intermediate balance between cosine-based and learned relevance provides the most effective interaction. A too-small $\alpha$ emphasizes the learned component and weakens geometric consistency, while an excessively large value diminishes adaptive relevance signals. Considering both performance and stability, we adopt $\alpha=0.3$ as the default setting in subsequent experiments.

### v2

@@ Abstract
Despite progress in multimodal sarcasm detection, existing datasets and methods predominantly focus on single-image scenarios, overlooking potential semantic and affective relations across multiple images. This leaves a gap in modeling cases where sarcasm is triggered by multi-image cues in real-world settings. To bridge this gap, we introduce MMSD3.0, a new benchmark composed entirely of multi-image samples curated from tweets and Amazon reviews. We further propose a Cross-Image Reasoning Model (CIRM), integrating a Dual-Stage Bridge Module and Relevance-Guided Fusion Module to model inter-image dependencies and cross-modal correspondences. Complementarily, we establish a comprehensive suite of strong and representative baselines and conduct extensive experiments, showing that MMSD3.0 is an effective and reliable benchmark that better reflects real-world conditions. Moreover, CIRM demonstrates state-of-the-art performance across MMSD, MMSD2.0, and MMSD3.0, validating its effectiveness in both single-image and multi-image scenarios. Dataset and code are publicly available at https://github.com/ZHCMOONWIND/MMSD3.0.
@@ 1 Introduction
Sarcasm is a common affective expression, yet it is hard to capture because the literal wording often contradicts the speaker’s true attitude, and it appears in diverse ironic forms. Accurately identifying sarcasm is therefore of substantial importance for emotion detection, online services, and opinion mining [36, 45, 21, 43, 48].
With the rapid surge of multimodal content on social media platforms, the expression of sarcasm has become increasingly diverse. In particular, sarcasm conveyed through both images and text poses significant challenges for automatic detection. This has led to the emergence of the multimodal sarcasm detection task. Cai et al. [3] were the first to construct the widely used MMSD dataset based on Twitter, providing a benchmark for research in this area. Early approaches [3, 43, 23, 41, 32], built on MMSD, fused visual and textual modalities to capture cross-modal incongruity and have achieved encouraging results. However, Qin et al. [33] later identified the presence of spurious cues and problematic samples within MMSD, and subsequently introduced an improved version, MMSD2.0. This new dataset addressed these issues and further advanced the development of more reliable multimodal sarcasm detection [37].
@@
A considerable fraction of tweets contain more than one image [^1] https://thenextweb.com/news/what-analyzing-1-million-tweets-taught-us, which is notably high. In such multi-image scenarios, sarcasm often arises from latent semantic or affective associations among the images. As shown in Figure 1, sarcasm can emerge from the contrast between images, for instance, the left image depicts Laura Loomer, while the right shows Count von Count. Without both images, the key contrast driving the sarcastic intent cannot be perceived. However, existing datasets [3, 33] and methods [43, 31, 23, 41, 32, 40, 4, 37, 30, 50] focus solely on single-image settings and cannot capture sarcasm driven by relations across multiple images. As a result, they do not faithfully reflect the complexity of multimodal sarcasm in real-world contexts.
@@
At the same time, to address the challenges of multi-image sarcasm detection, we propose CIRM. CIRM features a Dual-Stage Bridge Module (DSBM) and a Relevance-Guided Fusion Module (RGFM) for efficient cross-image and cross-modal reasoning. The bridge enables information exchange before and after sequence modeling, while the fusion leverages OCR-based alignment and adaptive relevance weights to integrate visual–textual cues. Image order is further encoded through positional embeddings and masking to enhance contextual understanding.
@@
  • We propose CIRM, a multi-image sarcasm reasoning framework that performs Dual-Stage Bridging and Relevance-Guided Fusion to enhance text–image alignment and improve multimodal understanding in complex sarcastic scenarios.
@@ 2.2 Multimodal Large Language Models
Table 1: Dataset statistics of MMSD, MMSD2.0, and MMSD3.0 across different splits.
Dataset Split Total Sar. Non-sar. Sar. (%) Avg. Len / Img
MMSD Train 19557 8385 11172 42.87 15.65 / 1.00
Val 2387 977 1410 40.93 15.68 / 1.00
Test 2373 969 1404 40.83 15.82 / 1.00
MMSD2.0 Train 19795 9557 10238 48.28 13.42 / 1.00
Val 2408 1040 1368 43.19 13.64 / 1.00
Test 2406 1034 1372 42.98 13.51 / 1.00
MMSD3.0 Train 7583 3007 4576 39.65 31.81 / 2.57
@@ 3.1 Data Collection
AI-generated Content and Augmentation. To mirror the surge of AI-generated deception on Twitter [35, 5, 10] and Amazon [44, 25, 28], we augment data using Qwen2.5-VL-32B [1] as the generator and GPT-4o as the evaluator. With two platform-specific prompts (Figures 8 and 9) in the Appendix, we input images from 1,444 real samples to generate three sarcastic candidates per sample. As shown in Appendix Figure 10, the evaluation prompt rates candidates on six criteria (e.g., Naturalness, Authenticity) and selects the top-scoring one, yielding 1,444 sarcasm-labeled samples.
@@ 3.2 Human Annotation
We recruited nine graduate students with backgrounds in multimodal research to conduct manual annotation. Given the inherently subjective nature of sarcasm [11, 46], each sample was assigned to two different annotators to reduce potential bias. Since understanding sarcasm in some cases may require specific background knowledge, annotators were allowed to search online for relevant context to better interpret the intended meaning. During the annotation process, we excluded any content containing violence, pornography, or other harmful or sensitive material. To evaluate annotation consistency, we used Cohen’s Kappa [27], and obtained a score of 0.816, indicating a high level of agreement [20]. Finally, we split the dataset into training, validation, and test sets following a 70:15:15 ratio.
@@ 4.4 Dual-Stage Bridge Module
Post-Bridge and Final Representation. After independent sequence modeling within each modality, we perform a second bridging stage to re-establish cross-modal alignment and enhance joint semantics for classification. The sequential block is first applied to both textual and visual streams:
@@
4.5 Relevance-Guided Fusion Module
@@
Relevance Estimation. A global textual summary $\bar{t}_{RGFM}$ is derived by averaging $T^{O}$. Each image representation $v_{i}^{O}$ is then scored against $\bar{t}$ from two perspectives: a cosine similarity term $s_{i}^{\cos}$ and a learnable term $s_{i}^{\mathrm{lrn}}$:
$s_{i}^{\cos}=\cos(W_{v}v_{i}^{O},\,W_{t}\bar{t}_{RGFM}),\qquad s_{i}^{\mathrm{lrn}}=\mathrm{MLP}\!\left([v_{i}^{O};\,\bar{t}]\right)$ (17)
@@ 4.6 Classification
After the Dual-Stage Bridge and Relevance-Guided Fusion Module, we obtain three complementary representations: the post-bridge text and vision streams $(T_{\text{post}},V_{\text{post}})$ and the relevance-guided feature $f$. We aggregate and integrate them in a shared fusion space for final prediction.
@@
where $c_{i}\in\{0,1\}$ masks padded images. The pooled features $\bar{t}$ and $\bar{v}$ are concatenated with the relevance-guided vector $f$ and an optional star-rating $r$ embedding:
$z=\mathrm{MLP}_{\text{fuse}}\!\left([\bar{t};\,\bar{v};\,f;\,\mathrm{Emb}(r)]\right)$ (21)
@@
$\mathcal{L}=\mathrm{CrossEntropy}(\hat{y},y;w)$ (23)
@@ 5 Experiments
Note. For details, please refer to Appendix §5.8 for the no-tiling protocol, §5.9 for real-world performance, and §5.11 for error cases.
@@ 5.1 Datasets
Because our data extraction differs from MMSD and MMSD2.0, we do not merge them. We first evaluated CIRM’s single-image performance on the two public datasets, MMSD and MMSD2.0. MMSD2.0 is an improved version of MMSD, with spurious cues removed and annotations refined, using training, validation, and test splits of 80%, 10%, and 10%, respectively. Subsequently, we built and evaluated multiple baselines on MMSD3.0 to assess performance in multi-image settings, with training, validation, and test splits of 70%, 15%, and 15% respectively.
@@ 5.2 Experimental Settings
We use Qwen2.5-VL-32B [1] as the generator and GPT-4o [29] as the evaluator. During baseline implementation, we reproduce each method following its official hyperparameters. CIRM is implemented based on the Hugging Face Transformers library and optimized using AdamW [26]. The learning rate is set to $2e^{-5}$ with a weight decay of $1e^{-5}$. The batch size is 8, and the model is trained for 20 epochs. All baseline construction and model training are conducted on a single NVIDIA H100 GPU (80 GB).
@@ 5.3 Baselines
Setting. We evaluate CIRM in both single-image and multi-image scenarios. For methods that only accept one image, we tile all $N$ images in order into a single canvas and keep the text unchanged, ensuring identical visual evidence without altering model designs.
Single-image. CIRM is evaluated against unimodal and multimodal baselines. Text-only: TextCNN [19] (CNN over $n$-grams); Bi-LSTM [12] (bidirectional RNN for long-range context); RoBERTa [24] (pretrained transformer encoder). Image-only: ResNet [14] (residual CNN); ViT [9] (Transformer over patch tokens). Multimodal: HFM [3] (hierarchical fusion of image, text, attributes); Att-BERT [31] (inter-modal attention on BERT); HKE [23] (graph reasoning over compositional congruity); DIP [41] (factual + affective cues mining); Multi-view CLIP [33] (diverse CLIP views for incongruity); Multi-view CLIP+TFCD [49] (training-free counterfactual debiasing); MoBA [42] (bi-directional LoRA with MoE routing); G2SAM [40] (graph-based global perception); CofiPara [4] (LLM-assisted sarcasm reasoning); Tang et al. [37] (instruction templates with demo retrieval); RCLMuFN [39] (relational learning with multiplex fusion).
Multi-image. CIRM and modern MLLMs natively accept multiple images; for single-image methods we use the tiling protocol above. Text-only: Bi-LSTM [12] (bidirectional recurrent backbone); BERT [8] (masked-language Transformer baseline); RoBERTa (TweetEval) [2] (robustly trained BERT variant for social text). Image-only: ResNet [14] (residual CNN on composite canvas); ViT [9] (patch-token Transformer on tiled images). Multimodal: DIP [41] (joint factual/affective fusion); Multi-view CLIP [33] (multi-perspective CLIP alignment); MoBA [42] (bidirectional modality interaction for robust fusion); Tang et al. [37] (LLM prompting with retrieved demonstrations). MLLMs: GPT-4o [29] (end-to-end multimodal reasoning with native multi-image input); LLaVA-1.5-7B [22] (vision-language fine-tuning on conversational data); Qwen2.5-VL-32B [1] (large-scale VL model with strong perception and instruction following). Prompts used for these MLLMs are provided in Appendix Figure 11.
@@ 5.5 Ablation Study
To verify the contribution of each component, we conduct ablation experiments under several settings: (1) w/o DSBM, removing the Dual-Stage Bridge Module; (2) w/o PE, removing positional encoding and masking; (3) w/o OCR, excluding OCR-based textual cues; (4) w/o Emoji, excluding emoji cues; (5) w/o RGFM, removing the Relevance-Guided Fusion Module; and (6) removing specific submodules within the Dual-Stage Bridge, including Pre-Bridge, Sequential Modeling, and Post-Bridge.
Table 4 summarizes the results. Removing the full DSBM or the Relevance-Guided Fusion Module causes the largest decline, highlighting their crucial roles in cross-modal reasoning. Excluding positional embeddings weakens performance, showing that image order matters for multi-image understanding. Omitting the Pre-Bridge or Post-Bridge slightly reduces accuracy, confirming their auxiliary value. The absence of OCR leads to a clear drop, as many images contain text crucial for sarcasm cues. Removing emoji features causes a small decline, indicating that emojis provide complementary affective signals. Overall, these results confirm that all modules jointly enhance multimodal reasoning on MMSD3.0.
Figure 5: Effect of mixing weight $\alpha$ on MMSD3.0, with the best F1-score ($84.42\%$) at $\alpha=0.3$.
5.6 Hyperparameter Study on Mixing Weight $\alpha$
We investigate the influence of the mixing weight $\alpha$ in the RGFM, where $s_{i}=\alpha\,s_{i}^{\mathrm{cos}}+(1-\alpha)\,s_{i}^{\mathrm{lrn}}$ balances cosine alignment with the learned scorer. We sweep $\alpha\in\{0,0.1,\ldots,1.0\}$ while keeping other settings fixed and evaluate on the MMSD3.0 test split. As shown in Figure 5, performance peaks at $\alpha=0.3$, achieving the highest F1-score of $84.42\%$. The accuracy remains stable around $84\%$ across most values, indicating robustness to moderate changes. Both very small and large $\alpha$ values, however, slightly reduce performance. These results suggest that a balance between cosine-based and learned relevance yields the best interaction. A small $\alpha$ emphasizes the learned component and weakens geometric consistency, while a large one diminishes adaptive relevance signals. Considering both performance and stability, we adopt $\alpha=0.3$ as the default in all experiments.
@@ 5.8 No-Tiling Multi-Image Experiment
Modality Method Real-world AI-Gen
Acc (%) F1 (%) Acc (%)
Multimodal DIP  [41] 79.59 75.50 97.98
Multi-view CLIP  [33] 80.01 75.48 95.96
MoBA  [42] 76.02 68.80 88.89
Tang et al.  [37] 80.36 75.93 95.45
MLLM GPT-4o  [29] 71.34 67.60 82.61
LLaVA-1.5-7B  [22] 59.62 55.78 72.83
Qwen2.5-VL-32B  [1] 68.90 66.65 95.65
Ours CIRM 83.31 80.39 98.48
@@
5.9 Real-World Performance
As illustrated in Table 6, there is a significant drop in model performance on real-world data compared to AI-generated samples, which underscores the inherent challenges posed by the greater complexity and unpredictability of real-world posts. The variation in context, language, and multimodal cues contributes to this discrepancy. Among the multimodal models, DIP and Tang et al. stand out with relatively competitive performance, whereas MoBA underperforms, and Multi-view CLIP exhibits limited adaptability to real-world conditions. On the MLLM front, while GPT-4o and Qwen2.5-VL-32B deliver moderate accuracy, LLaVA-1.5-7B lags behind in both accuracy and F1 score.
In contrast, CIRM consistently outperforms all baselines, achieving a robust F1 score of $80.39\%$ on real-world data and a remarkable $98.48\%$ on AI-generated samples. This reinforces the model’s versatility and highlights the effectiveness of its cross-image reasoning and Relevance-Guided Fusion strategies. By focusing on cross-modal consistency and latent inter-image relationships, CIRM demonstrates superior capacity to capture and interpret the subtle, multifaceted nature of sarcasm in real-world multimodal data. The results not only confirm the importance of tailored cross-image reasoning but also establish CIRM as a promising solution for multimodal sarcasm detection in complex, varied environments.